### v1

# Non-isogenous superelliptic jacobians II

Abstract.
Let $\ell$ be an odd prime and $K$ a field of characteristic different from $\ell$. Let $\bar{K}$ be an algebraic closure of $K$. Assume that $K$ contains a primitive $\ell$th root of unity. Let $n\neq\ell$ be another odd prime. Let $f(x)$ and $h(x)$ be degree $n$ polynomials with coefficients in $K$ and without repeated roots.
Let us consider superelliptic curves $C_{f,\ell}:y^{\ell}=f(x)$ and $C_{h,\ell}:y^{\ell}=h(x)$ of genus $(n-1)(\ell-1)/2$, and their jacobians $J^{(f,\ell)}$ and $J^{(h,\ell)}$, which are $(n-1)(\ell-1)/2$-dimensional abelian varieties over $\bar{K}$.
Suppose that one of the polynomials is irreducible and the other reducible over $K$. We prove that if $J^{(f,\ell)}$ and $J^{(h,\ell)}$ are isogenous over $\bar{K}$ then both endomorphism algebras $\mathrm{End}^{0}(J^{(f,\ell)})$ and $\mathrm{End}^{0}(J^{(h,\ell)})$ contain an invertible element of multiplicative order $n$.
Key words and phrases:
superelliptic curves, jacobians, isogenies of abelian varieties
2010 Mathematics Subject Classification:
14H40, 14K05, 11G30, 11G10
The author was partially supported by Simons Foundation Collaboration grant # 585711. This work was done during his stay in 2022 and 2023 at the Max-Planck Institut für Mathematik (Bonn, Germany), whose hospitality and support are gratefully acknowledged.
1. Definitions, notations, statements
This paper is a follow up of [18, 21] and we use its (more or less standard) notation. (See also [16, 20, 21].) In particular, $\ell$ is an odd prime, ${\mathbb{F}}_{\ell}$ the corresponding (finite) prime field of characteristic $\ell$. We write ${\mathbb{Z}}_{\ell}$ and ${\mathbb{Q}}_{\ell}$ for the ring of $\ell$-adic integers and the field ${\mathbb{Q}}_{\ell}$ of $\ell$-adic numbers respectively. Let us fix a primitive $\ell$th root of unity
$\zeta_{\ell}\in{\mathbb{C}}$
in the field ${\mathbb{C}}$ of complex numbers. We write ${\mathbb{Q}}(\zeta_{\ell})$ for the $\ell$th cyclotomic field and
${\mathbb{Z}}[\zeta_{q}]=\sum_{i=0}^{\phi(q)-1}{\mathbb{Z}}\cdot\zeta_{q}^{i}$
for its ring of integers. We write $\mathcal{P}_{\ell}(t)$ for the $\ell$th cyclotomic polynomial
$\mathcal{P}_{\ell}(t)=\sum_{j=0}^{\ell-1}t^{j}\in{\mathbb{Z}}[t].$
Let $K$ be a field with $\mathrm{char}(K)\neq\ell$. Let us fix an algebraic closure $\bar{K}$ of $K$ and write $\mathrm{Gal}(K)=\mathrm{Aut}(\bar{K}/K)$ for its group of its $K$-linear automorpisms. In what follows we always assume that $K$ contains a primitive $\ell$th root of unity, say, $\zeta$.
Let $f(x)\in K[x]$ be a polynomial with coefficients in $K$, of degree $n$ not diivisible by $\ell$, and without repeated roots. We write $\mathfrak{R}_{f}\subset\bar{K}$ for the $n$-element set of roots of $f(x)$, $K(\mathfrak{R}_{f})$ for the splitting field of $f(x)$ and $\mathrm{Gal}(f/K)$ for the Galois group
$\mathrm{Gal}(K(\mathfrak{R}_{f})/K)=\mathrm{Aut}(K(\mathfrak{R}_{f})/K)$
of $f(x)$. As usual, one may view $\mathrm{Gal}(f/K)$ as a certain permutation subgroup of the group $\mathrm{Perm}(\mathfrak{R}_{f})$ of all permutations of $\mathfrak{R}_{f}$.
We write $C_{f,\ell}$ for the smooth projective model of the plane affine curve $y^{\ell}=f(x)$. It is well known [10, 11] that the genus $g(C_{f,\ell})$ of $C_{f,\ell}$ is $(\ell-1)(n-1)/2$ if $\ell$ does not divide $n$. Further we assume that $\ell\nmid n$. The map
$(x,y)\mapsto(x,\zeta y)$
gives rise to a non-trivial biregular automorphism
$\delta_{\ell}:C_{f,\ell}\to C_{f,\ell}$
of period $\ell$ that is defined over $K$.
Let $J^{(f,\ell)}$ be the jacobian of $C_{f,\ell}$; it is a $(\ell-1)(n-1)/2$-dimensional abelian variety that is defined over $K$. We write $\mathrm{End}(J^{(f,\ell)})$ for the ring of $\bar{K}$-endomorphisms of $J^{(f,\ell)}$ and $\mathrm{End}^{0}((J^{(f,\ell)})=\mathrm{End}(J^{(f,\ell)})\otimes{\mathbb{Q}}$ for the corresponding endomorphism algebra of $J^{(f,\ell)}$. By functoriality, $\delta_{\ell}$ induces a $K$-automorphism of $J^{(f,\ell)}$, which we still denote by $\delta_{q}$. It is known ([10, p. 149], [11, p. 448]; see also [19]) that
$\mathcal{P}_{\ell}(\delta_{\ell})=\sum_{j=0}^{\ell-1}\delta_{\ell}^{j}=0$ (1)
in $\mathrm{End}(J^{(f,\ell)})$. Then (1) gives rise to the ring homomorphism,
${\bf i}_{\ell,f}:{\mathbb{Z}}[\zeta_{\ell}]\hookrightarrow{\mathbb{Z}}[\delta_{\ell}]\subset\mathrm{End}(J^{(f,\ell)}),\ \zeta_{\ell}\mapsto\delta_{\ell},$ (2)
which is a ring embedding ([10, p. 149], [11, p. 448]; see also [19]). This implies that the subring ${\mathbb{Z}}[\delta_{\ell}]$ of $\mathrm{End}(J^{(f,\ell)})$ generated by $\delta_{\ell}$ is isomorphic to ${\mathbb{Z}}[\zeta_{\ell}]$. It follows that the ${\mathbb{Q}}$-subalgebra
${\mathbb{Q}}[\delta_{\ell}]\subset\mathrm{End}^{0}(J^{(f,\ell)})$ (3)
generated by $\delta_{\ell}$ is canonically isomorphic to the $\ell$th cyclotomic field ${\mathbb{Q}}(\zeta_{\ell})$ and therefore has ${\mathbb{Q}}$-dimension $\ell-1$.
Throughout this paper, $\ell$ is a prime, $n\geq 3$ is an integer not divisible by $\ell$, and $f(x)$ and $h(x)$ are degree $n$ polynomials with coefficients in $K$ and without repeated roots,
$C_{f,\ell}:y^{\ell}=f(x),\ C_{h,\ell}:y^{\ell}=h(x)$
are the corresponding genus $(n-1)(\ell-1)/2$ superelliptic curves over $K$, whose jacobians we denote by $J^{(f,\ell)}$ and $J^{(h,\ell)}$, respectively. These jacobians are $(n-1)(\ell-1)/2$-dimensional abelian varieties defined over $K$. Hereafter $\ell$ is an odd prime that does not divide $n$.
The main result of this paper is the following assertion.
Theorem 1.1.
Suppose that $n$ and $\ell$ are distinct odd primes. Let $K$ be a field of characteristic different from $\ell$. Let $f(x),h(x)\in K[x]$ be degree $n$ polynomials without repeated roots. Suppose that one of the polynomials is irreducible and the other is reducible.
Then both endomorphism algebras $\mathrm{End}^{0}(J^{(f,\ell)})$ and $\mathrm{End}^{0}(J^{(h,\ell)})$ contain an invertible element of multiplicative order $n$.
The next assertion may be viewed as a partial generalization of Theorem 1.1 to the case of arbitrary odd $n$.
Theorem 1.2.
Suppose that $n\geq 3$ is an odd integer and $\ell$ is an odd prime not dividing $n$. Let $K$ be a field of characteristic different from $\ell$. Let $f(x),h(x)\in K[x]$ be degree $n$ polynomials without repeated roots. Suppose that $f(x)$ is irreducible over $K$.
Assume additionally that the order of the Galois group $\mathrm{Gal}(h/K)$ of $h(x)$ is prime to $n$. (E.g., each irreducible factor of $h(x)$ over $K$ has degree $1$ or $2$.)
If the corresponding superelliptic jacobians $J^{(f,\ell)}$ and $J^{(h,\ell)}$ are isogenous over $\bar{K}$ then there is a prime divisor $r$ of $n$ such that both endomorphism algebras $\mathrm{End}^{0}(J^{(f,\ell)})$ and $\mathrm{End}^{0}(J^{(h,\ell)})$ contain an invertible element of multiplicative order $r$.
Remark 1.3.
If the conditions of Theorem 1.2 hold then $h(x)$ is reducible over $K$, see [21, Remark 1.4].
Example 1.4.
Let $\ell$ be an odd prime and $n\geq 3$ an odd integer that is not divisible by $\ell$. Let us take as $K$ the $\ell$th cyclotomic field ${\mathbb{Q}}(\zeta_{\ell})$. Let us put
$f(x)=x^{n}-2,\ h(x)=x^{n}-1\in K[x].$
By Eisenstein criterion, $f(x)$ is irreducible over $K$ (recall a prime $2$ is unramified in ${\mathbb{Q}}(\zeta_{\ell})=K$, because $\ell$ is odd) while $h(x)$ is obviously reducible over $K$. Let $\ell$ be an odd prime that does not divide $n$. The curves $C_{f,\ell}$ and $C_{h,\ell}$ are obviously isomorphic over $\bar{K}$. They both admit periodic $\bar{K}$-automorphisms $\tilde{\delta}_{n}$ of order $n$ defined by the (same) formula
$(x,y)\mapsto(\zeta_{n}x,y)$
where
$\zeta_{n}\in\bar{K}=\bar{{\mathbb{Q}}}\subset{\mathbb{C}}$
is a primitive $n$th root of unity. This implies that their jacobians $J^{(f,\ell)}$ and $J^{(h,\ell)}$ are abelian varieties over $\bar{K}$ that are isomorphic over $\bar{K}$ and admit periodic automorphisms of (multiplicative) order $n$ that we continue to denote by $\tilde{\delta}_{n}$. It follows that if $r$ is any prime divisor of $n$ then both $J^{(f,\ell)}$ and $J^{(h,\ell)}$ admit automorphisms $\left(\tilde{\delta}_{n}\right)^{n/r}$ of multiplicative order $r$ that may be viewed as invertible elements of multiplicative order $r$ in $\mathrm{End}^{0}(J^{(f,\ell)})$ and $\mathrm{End}^{0}(J^{(h,\ell)})$.
Example 1.5.
Let $n\geq 5$ be an odd integer and $\ell$ an odd prime not dividing $n$. Let us put $K={\mathbb{Q}}(\zeta_{n})$.
Let $f(x)\in K[x]$ be a degree $n$ irreducible polynomial over $K$, whose Galois group $\mathrm{Gal}(f/K)$ is either the full symmetric group $\mathrm{S}_{n}$ or the alternating group $\mathrm{A}_{n}$. It is known [15, 17, 22] that the endomorphism algebra $\mathrm{End}^{0}(J^{(f,\ell)})$ is isomorphic to the field ${\mathbb{Q}}(\zeta_{\ell})$; in particular, $J^{(f,\ell)}$ is absolutely simple. The (cyclic) multiplicative group of all roots of unity in ${\mathbb{Q}}(\zeta_{\ell})$ has order $2\ell$, which is prime to (odd) $n$. Hence, $\mathrm{End}^{0}(J^{(f,\ell)})\cong{\mathbb{Q}}(\zeta_{\ell})$ does not contain elements of multiplicative order $r$ for any prime divisor $r$ of $n$.
Let $h(x)\in K[x]$ be a degree $n$ polynomial that splits over $K$ into a product of linear factors.
Then it follows from Theorem 1.2 that $J^{(f,\ell)}$ and $J^{(h,\ell)}$ are not isogenous over $\bar{K}=\bar{{\mathbb{Q}}}$ (and therefore even over $\bar{{\mathbb{C}}}$). Since $J^{(f,\ell)}$ is absolutely simple and has the same dimension as $J^{(h,\ell)}$, it follows that every ${\mathbb{C}}$-homomorphism between $J^{(f,\ell)}$ and $J^{(h,\ell)}$ is zero.
Example 1.6.
Let $n\geq 5$ be an odd integer and $\ell$ an odd prime not dividing $n$. Let us consider the degree $n$ polynomials
$f_{1}(x)=x^{n}-x-1,\quad f_{2}(x)=\sum_{j=0}^{n}x^{j}/j!$
with rational coefficients. It is known (Selmer, Osada [8]) that $f_{1}(x)$ is irreducible over ${\mathbb{Q}}$ and $\mathrm{Gal}(f_{1}/{\mathbb{Q}})=\mathrm{S}_{n}$. By a theorem of Schur [1], $f_{2}(x)$ is irreducible over ${\mathbb{Q}}$ and $\mathrm{Gal}(f_{2}/{\mathbb{Q}})=\mathrm{S}_{n}$ or $\mathrm{A}_{n}$.
Recall that $n\geq 5$ and therefore $\mathrm{A}_{n}$ is a simple non-abelian group that coincides with the commutator subgroup of $\mathrm{S}_{n}$; in addition, $\mathrm{A}_{n}$ is a maximal subgroup of $\mathrm{S}_{n}$. Since $K={\mathbb{Q}}(\zeta_{\ell})$ is an abelian extension of ${\mathbb{Q}}$, the Galois groups $\mathrm{Gal}(f_{1}/K)$ and $\mathrm{Gal}(f_{2}/K)$ are either $\mathrm{S}_{n}$ or $\mathrm{A}_{n}$; in particular, both $f_{1}$ and $f_{2}$ remain irreducible over $K$.
It follows from Example 1.5 that if $h(x)\in K[x]$ is a degree $n$ polynomial that splits over $K$ into a product of linear factors then every ${\mathbb{C}}$-homomorphism between $J^{(f_{k},\ell)}$ and $J^{(h,\ell)}$ is zero (for both $k=1,2$).
The paper is organized as follows. In Section 2 we recall basic facts about Galois properties of points of order $\ell$ on superelliptic jacobians. We also state Theorem 2.2 that is a slightly stronger version of Theorem 1.2. Section 3 contains the proof of Theorem 2.2. We prove Theorem 1.1 in Section 4.
2. Points of order $\ell$ on superelliptic jacobians
Let $K_{s}\subset\bar{K}$ be the separable algebraic closure of $K$. Let $X$ be a positive-dimensional abelian variety over $K$. If $d$ is a positive integer then we write $X[d]$ for the kernel of multiplication by $d$ in $X(\bar{K})$. Recall ([7, Sect. 6], [5, Sect. 8, Remark 8.4]) that if $d$ is not divisible by $\mathrm{char}(K)$ then $X[d]$ is a $\mathrm{Gal}(K)$-submodule of $X(K_{s})$; in addition, $X[d]$ is isomorphic as a commutative group to $({\mathbb{Z}}/d{\mathbb{Z}})^{2\mathrm{dim}(X)}$.
Let $K(X[d])$ be the field of definition of all torsion points of order dividing $d$ on $X$. It is well known [5, Remark 8.4] that $K(X[d])$ lies in $K_{s}$ and is a finite Galois extension of $K$. Let us put
$\tilde{G}_{d,X}:=\mathrm{Gal}(K(X[d])/K).$
One may view $\tilde{G}_{d,X}$ as a certain subgroup of $\mathrm{Aut}_{{\mathbb{Z}}/d{\mathbb{Z}}}(X[d])$ and $X[d]$ as a faithful $\tilde{G}_{d,X}$-module. In addition, the structure of the $\mathrm{Gal}(K)$-module on $X[d]$ is induced by the canonical (continuous) surjective group homomorphism
$\tilde{\rho}_{d,X}:\mathrm{Gal}(K)\twoheadrightarrow\mathrm{Gal}(K(X[d])/K)=\tilde{G}_{d,X}.$
For example, if $d$ is a prime $\ell\neq\mathrm{char}(K)$ then $X[\ell]$ is a $2\mathrm{dim}(X)$-dimensional vector space over the $\ell$-element prime field ${\mathbb{F}}_{\ell}={\mathbb{Z}}/\ell{\mathbb{Z}}$ and the inclusion $\tilde{G}_{\ell,X}\subset\mathrm{Aut}_{{\mathbb{F}}_{\ell}}(X[\ell])$ defines a faithful linear representation of the group $\tilde{G}_{\ell,X}$ in the vector space $X[\ell]$ over ${\mathbb{F}}_{\ell}$.
Remark 2.1.
If $d\geq 3$ is an integer that is every ${\mathbb{C}}$-homomorphism between $J^{(f,\ell)}$ and $J^{(h,\ell)}$ is zero.not divisible by $\mathrm{char}(K)$ then all the endomorphisms of $X$ are defined over $K(X[d])$, by a theorem of A. Silverberg [14, Th. 2.4]. (See also [3, 12, 2].)
The following assertion may be viewed as a variant of Theorem 1.2 (when $Y=J^{(h,\ell)}$).
Theorem 2.2.
Let $\ell$ be an odd prime and $K$ a field of characteristic $\neq\ell$. Let $n\geq 3$ be an odd positive integer that is not divisible by $\ell$, and $f(x)\in K[x]$ a degree $n$ irreducible polynomial without repeated roots. Let us put $X=J^{(f,\ell)}$, which is a $(n-1)(\ell-1)/2$-dimensional abelian variety over $K$.
Let $Y$ be a $(n-1)(\ell-1)/2$-dimensional abelian variety over $K$ such that the order of $\tilde{G}_{\ell,Y}$ is prime to $n$. (E.g., $K(Y[\ell])=K$ or this order is a power of $\ell$.)
Suppose that $X$ and $Y$ are isogenous over $\bar{K}$.
Then there is an odd prime $r$ dividing $n$ such that both endomorphism algebras $\mathrm{End}^{0}(X)$ and $\mathrm{End}^{0}(Y)$ contain an invertible element of multiplicative order $r$.
We will prove Theorem 2.2 in Section 3. Our proof is based on the Galois properties of certain points of order $\ell$ on superelliptic jacobians $J^{(f,\ell)}$ that will be discussed in the next subsection.
2.1. Galois properties
In this subsection we recall an explicit description of a certain important Galois submodule of $J^{(f,\ell)}[\ell]$ [10, 11] for arbitrary separable $f(x)$ and odd $n$ under an additional assumption that $\ell$ does not divide $n$.
Let us start with the $n$-dimensional ${\mathbb{F}}_{\ell}$-vector space
${\mathbb{F}}_{\ell}^{\mathfrak{R}_{f}}=\{\phi:\mathfrak{R}_{f}\to{\mathbb{F}}_{\ell}\}$
of all ${\mathbb{F}}_{\ell}$-valued functions on $\mathfrak{R}_{f}$. The action of $\mathrm{Perm}(\mathfrak{R}_{f})$ on $\mathfrak{R}_{f}$ provides ${\mathbb{F}}_{\ell}^{\mathfrak{R}_{f}}$ with the structure of a faithful $\mathrm{Perm}(\mathfrak{R}_{f})$-module, which splits into a direct sum
${\mathbb{F}}_{\ell}^{\mathfrak{R}_{f}}={\mathbb{F}}_{\ell}\cdot{\bf 1}_{\mathfrak{R}_{f}}\oplus Q_{\mathfrak{R}_{f}}$ (4)
of the one-dimensional subspace ${\mathbb{F}}_{\ell}\cdot{\bf 1}_{\mathfrak{R}_{f}}$ of constant functions and the $(n-1)$-dimensional heart [4, 6]
$Q_{\mathfrak{R}_{f}}:=\{\phi:\mathfrak{R}_{f}\to{\mathbb{F}}_{\ell}\mid\sum_{\alpha\in\mathfrak{R}_{f}}\phi(\alpha)=0\}$
(here we use that $n$ is not divisible by $\ell$). Clearly, the $\mathrm{Perm}(\mathfrak{R}_{f})$-module $Q_{\mathfrak{R}_{f}}$ is faithful. It remains faithful if we view it as a $\mathrm{Gal}(f/K)$-module.
Remark 2.3.
There is a nondegenerate $\mathrm{Perm}(\mathfrak{R}_{f})$-invariant ${\mathbb{F}}_{\ell}$-bilinear pairing
$\Psi:{\mathbb{F}}_{\ell}^{\mathfrak{R}_{f}}\times{\mathbb{F}}_{\ell}^{\mathfrak{R}_{f}}\to{\mathbb{F}}_{\ell},\ \phi,\psi\mapsto\sum_{\alpha\in\mathfrak{R}_{f}}\phi(\alpha)\psi(\alpha)$
and the splitting (4) is an orthogonal direct sum. Clearly, the restriction of $\Psi$ to ${\mathbb{F}}_{2}\cdot{\bf 1}_{\mathfrak{R}_{f}}$ is nondegenerate and therefore the restriction of $\Psi$ to $Q_{\mathfrak{R}_{f}}$ is nondegenerate as well. This implies that the $\mathrm{Gal}(f/K)$-module $Q_{\mathfrak{R}_{f}}$ and its dual $\mathrm{Hom}_{{\mathbb{F}}_{\ell}}(Q_{\mathfrak{R}_{f}},{\mathbb{F}}_{\ell})$ are isomorphic.
The field inclusion $K(\mathfrak{R}_{f})\subset K_{s}$ induces the surjective continuous group homomorphism
$\mathrm{Gal}(K)=\mathrm{Gal}(K_{s}/K)\twoheadrightarrow\mathrm{Gal}(K(\mathfrak{R}_{f})/K)=\mathrm{Gal}(f/K),$
which gives rise to the natural structure of the $\mathrm{Gal}(K)$-module on $Q_{\mathfrak{R}_{f}}$ such that the image of $\mathrm{Gal}(K)$ in $\mathrm{Aut}_{{\mathbb{F}}_{\ell}}(Q_{\mathfrak{R}_{f}})$ coincides with
$\mathrm{Gal}(f/K)\subset\mathrm{Perm}(\mathfrak{R}_{f})\hookrightarrow\mathrm{Aut}_{{\mathbb{F}}_{\ell}}(Q_{\mathfrak{R}_{f}}).$
In order to explain why the structure of the Galois module $Q_{\mathfrak{R}_{f}}$ is important, let us consider the subgroup (actually, the Galois submodule)
$J^{(f,\ell)}[1-\delta_{\ell}]=\{z\in J^{(f,\ell)}(\bar{K})\mid\delta_{l}(z)=z\}$
of $J^{(f,\ell)}(\bar{K})$. B. Poonen and E. Schaefer [10, 11] observed that the Galois module $J^{(f,\ell)}[1-\delta_{\ell}]$ is a Galois submodule of $J^{(f,\ell)}[\ell]$ and is isomorphic to $Q_{\mathfrak{R}_{f}}$. In particular, $K(\mathfrak{R}_{f})$ coincides with the field of definition of all points of $J^{(f,\ell)}[1-\delta_{\ell}]$.
We will need the following elementary assertion about homomorphisms of Galois modules related to $Q_{\mathfrak{R}_{f}}$.
Lemma 2.4.
Suppose that $f(x)$ is irreducible over $K$ and a prime $\ell$ does not divide $n$. Then:
  • (i)
    $Q_{\mathfrak{R}_{f}}$ does not contain nonzero Galois-invariants.
  • (ii)
    Every Galois-invariant linear functional $Q_{\mathfrak{R}_{f}}\to{\mathbb{F}}_{\ell}$ is zero.
  • (iii)
    Let $W$ be a ${\mathbb{F}}_{\ell}$-vector space provided with the trivial action of $\mathrm{Gal}(K)$. Then every homomorphism of Galois modules
    $Q_{\mathfrak{R}_{f}}\to W$
    is zero.
  • (iv)
    Let $V$ be a finite-dimensional ${\mathbb{F}}_{\ell}$-vector space provided with a linear action of $\mathrm{Gal}(K)$ in such a way that every simple (Jordan-Hölder) subquotient of $V$ is a trivial Galois module. Then every homomorphism of the Galois modules $Q_{\mathfrak{R}_{f}}\to V$ is zero.
  • (v)
    Let $V$ be a finite-dimensional ${\mathbb{F}}_{\ell}$-vector space provided with a linear action of $\mathrm{Gal}(K)$ in such a way that every simple (Jordan-Hölder) subquotient of $V$ is a trivial Galois module. Let $M$ be a finite-dimensional ${\mathbb{F}}_{\ell}$-vector space provided with a linear action of $\mathrm{Gal}(K)$ in such a way that there is a filtration
    $M_{0}=\{0\}\subset M_{1}\subset\dots\subset M_{d}=M$
    of $M$ by Galois submodules $M_{i}$ such that every quotient $M_{i+1}/M_{i}$ is isomorphic to $Q_{\mathfrak{R}_{f}}$. Then every homomorphism of the Galois modules $M\to V$ is zero.
Proof.
Recall that the irreducibility means that the Galois group acts transitively on $\mathfrak{R}_{f}$. Let $\phi\in Q_{\mathfrak{R}_{f}}$ be a Galois-invariant function on $\mathfrak{R}_{f}$. The transitivity implies that $\phi$ is constant. This means that there is $c\in{\mathbb{F}}_{\ell}$ such that $\phi(\alpha)=c$ for all $\alpha\in\mathfrak{R}_{f}$ and therefore (since $\phi\in Q_{\mathfrak{R}_{f}}$)
$0=\sum_{\alpha\in\mathfrak{R}_{f}}\phi(\alpha)=n\cdot c,$
i.e., $c=0$. This means that $\phi\equiv 0$, which proves (i). In order to prove the second assertion of Lemma, recall (Remark 2.3) that the Galois modules $Q_{\mathfrak{R}_{f}}$ and $\mathrm{Hom}_{{\mathbb{F}}_{\ell}}(Q_{\mathfrak{R}_{f}},{\mathbb{F}}_{\ell})$ are isomorphic. Now the second assertion of our Lemma follows from the already proven first one. On the other hand, the third assertion is an immmediate corollary of the second one: one has only to choose a basis of $W$. In order to prove (iv), we will use induction by $\mathrm{dim}(V)$. We may assume that the subspace $V\neq\{0\}$. It follows from our assumptions on the Galois module $V$ that the subspace $V_{0}=V^{\mathrm{Gal}(K)}$ of Galois invariants is not $\{0\}$ as well. If $u:Q_{\mathfrak{R}_{f}}\to V$ is a homomorphism of Galois modules then the induction assumption applied to the quotient $V/V_{0}$ implies that the induced homomorphism of Galois modules
$Q_{\mathfrak{R}_{f}}\to V/V_{0},\ \phi\mapsto u(\phi)+V_{0}$
is zero. This means that $u(Q_{\mathfrak{R}_{f}})\subset V$. Now the desired result follows from (iii) applied to
$Q_{\mathfrak{R}_{f}}\to V_{0},\ \phi\mapsto u(\phi)\in V_{0},$
because $\mathrm{Gal}(K)$ acts trivially on $V_{0}$.
In order to prove (v), let us use induction by $d$. Let $u:M\to V$ be a homomorphism of Galois modules. Since $M_{1}$ is isomorphic to $Q_{\mathfrak{R}_{f}}$, it follows from (iv) that $u(M_{1})=\{0\}$, i.e., there is a homomorphism of Galois modules $u_{1}:M/M_{1}\to V$ such that $u$ is the composition of
$M\to M/M_{1},\ m\mapsto m+M_{1}$
and $u_{1}$. If $d=1$ then we are done. If $d>1$ then the desired result follows from the induction assumption applied to the filtered Galois module
$M/M_{1}=(M/M_{1})_{d-1}\supset\dots\supset M_{1}/M_{1}=\{0\}=(M/M_{0})_{0}.$
3. Isogenous superelliptic jacobians
We will deduce Theorem 2.2 from the following auxiliary statements.
Lemma 3.1 (See Lemma 3.1 of [21]).
Let $G$ be a transitive permutation group of a finite nonempty set $\mathfrak{R}$, and $H$ a normal subgroup of $G$. Then the number of $H$-orbits in $\mathfrak{R}$ divides both $\#(\mathfrak{R})$ and the index $(G:H)$. In particular, if $\#(\mathfrak{R})$ and $(G:H)$ are relatively prime then $H$ acts transitively on $\mathfrak{R}$.
Lemma 3.2 (See Lemma 3.2 of [21]).
Let $f(x)$ be a degree $n$ irreducible polynomial over a field $K$ and without repeated roots. Let $K_{1}/K$ be a finite Galois field extension, whose degree is prime to $n$. Then $f(x)$ remains irreducible over $K_{1}$. In particular, the order of Galois group $\mathrm{Gal}(f/K_{1})$ is divisible by $n$.
Lemma 3.3.
Let $\ell$ be a prime, $F$ a field of prime characteristic $\ell$, and $V$ a finite-dimensional vector space over $F$. Let $\mathcal{N}\subset\mathrm{End}_{F}(V)$ be a linear nilpotent Lie subalgebra of $\mathrm{End}_{F}(V)$, and
$V^{\mathcal{N}}:=\{v\in V\mid x(v)=0\ \forall x\in\mathcal{N}\}.$
Let $\sigma$ be a linear automorphism of finite order in $V$ that commutes with all linear operators from $\mathcal{N}$ and such that the subspace $V^{\sigma}$ of all $\sigma$-invariants in $V$ contains $V^{\mathcal{N}}$. Then the order of $\sigma$ is either $1$ or a power of $\ell$.
Remark 3.4.
We will apply the easy “commutative” case of Lemma 3.3 to $F={\mathbb{F}}_{\ell}$, $V=X[\ell]$, and
$\sigma\in\mathrm{Gal}(K(X[\ell]/K(X[1-\delta_{\ell}]))\subset\mathrm{Aut}_{{\mathbb{F}}_{\ell}}(X[\ell])$
where $X=J^{(f,\ell)}$.
Proof of Lemma 3.3.
Replacing $\sigma$ by its suitable power, we may assume that $\sigma$ is a periodic automorphism, whose order is prime to $\ell$. We need to prove that $\sigma$ is the identity map.
Since $\sigma$ is obviously a semisimple linear operator, $V$ splits into a direct sum of $\sigma$-invariant subspaces
$V=V^{\sigma}\oplus W\ \text{ where }W=(1-\sigma)V.$
In particular,
$V^{\sigma}\cap W=\{0\}.$
Assume that $W\neq\{0\}$. We need to arrive to a contradiction. Since $\mathcal{N}$ commutes with $\sigma$, the subspace $W=(1-\sigma)V$ is $\mathcal{N}$-invariant. By Engel’s theorem, there is a nonzero vector $w\in W$ that is killed all linear operators from $\mathcal{N}$, i.e., $w\in V^{\mathcal{N}}$. This implies that $w\in W\in V^{\sigma}$. Hence,
$w\in V^{\sigma}\cap W=\{0\}.$
This implies that $w=0$, which gives us a desired contradiction.
Corollary 3.5.
Let $X=J^{(f,\ell)}$. Then:
  • (i)
    $K(X[\ell])/K(\mathfrak{R}_{f})$ is a finite Galois $\ell$-extension, i.e., either $K(X[\ell])=K(\mathfrak{R}_{f})$ or the Galois group $\mathrm{Gal}(K(X[\ell])/K(\mathfrak{R}_{f}))$ is a finite $\ell$-group.
  • (ii)
    The Galois module $X[\ell]$ admits a filtration
    $M_{0}=\{0\}\subset M_{1}\subset\dots\subset M_{\ell-1}=X[\ell],$
    such that each consecutive quotient $M_{i+1}/M_{i}$ is isomorphic to the Galois module $Q_{\mathfrak{R}_{f}}$.
Proof.
Let us put
$X[1-\delta_{\ell}]:=J^{(f,\ell)}[1-\delta_{\ell}]\subset J^{(f,\ell)}[\ell]=X[\ell].$
The ring ${\mathbb{Z}}[\delta_{\ell}]\otimes{\mathbb{Z}}_{\ell}=:{\mathbb{Z}}_{\ell}[\delta_{\ell}]$ acts naturally on the $\ell$-adic Tate module $\mathrm{T}_{\ell}(X)$ of $X$. It is known [13] that $\mathrm{T}_{\ell}(X)$ is a free ${\mathbb{Z}}_{\ell}[\zeta_{\ell}]$-module of rank
$\frac{2\mathrm{dim}(J^{(f,\ell)})}{[{\mathbb{Q}}(\zeta_{\ell}):{\mathbb{Q}}]}=n-1.$
In particular, the natural ring homomorphism
$\Psi_{\ell}:{\mathbb{Z}}[\delta_{\ell}]/\ell={\mathbb{Z}}_{\ell}[\delta_{\ell}]/\ell\to\mathrm{End}_{{\mathbb{F}}_{\ell}}(X[\ell])$
is a ring embedding that makes $X[\ell]$ a free ${\mathbb{Z}}[\delta_{\ell}]/\ell$-module. This implies that
$X[1-\delta_{\ell}]=(1-\delta_{\ell})^{\ell-2}X[\ell],$
because in the cyclotimic ring ${\mathbb{Z}}[\zeta_{\ell}]$ we have the equalities of ideals
$\ell{\mathbb{Z}}[\zeta_{\ell}]=(1-\zeta_{\ell})^{\ell-1}{\mathbb{Z}}[\zeta_{\ell}],\quad(1-\zeta_{\ell})^{\ell-2}{\mathbb{Z}}[\zeta_{\ell}]=\{z\in{\mathbb{Z}}[\zeta_{\ell}]\mid(1-\zeta_{\ell})z\in\ell{\mathbb{Z}}[\zeta_{\ell}]\}.$
Now (i) follows from Lemma 3.3 applied to
$F={\mathbb{F}}_{\ell},\ V=X[\ell],\ \mathcal{N}=\Psi_{\ell}((1-\delta_{\ell}){\mathbb{Z}}[\delta_{\ell}]/\ell),$
and
$\sigma\in\mathrm{Gal}(K(X[\ell]/K(X[1-\delta_{\ell}))\subset\mathrm{Aut}_{{\mathbb{F}}_{\ell}}(X[\ell]).$
In order to prove (ii), let us put
$M_{i}:=(1-\delta_{\ell})^{\ell-1-i}X[\ell]\subset X[\ell].$
The freeness of the ${\mathbb{Z}}[\delta_{\ell}]/\ell$-module $X[\ell]$ implies that $M_{i}$ coincides with the kernel of
$(1-\delta_{\ell})^{i}:X[\ell]\to X[\ell].$
In particular,
$M_{0}=\{0\},\ M_{1}=X[1-\delta_{\ell}]\cong Q_{\mathfrak{R}_{f}},\ M_{\ell-1}=X[\ell].$
It is also clear that $(1-\delta_{\ell})^{i}$ induces an isomorphism of Galois modules $M_{i+1}/M_{i}$ and $M_{1}/M_{0}\cong Q_{\mathfrak{R}_{f}}$. This ends the proof of (ii).
Lemma 3.6.
We keep the notation and assumptions of Theorem 2.2.
Suppose that $K=K(Y[\ell])$. Then there is a nontrivial group homomorphism
$\chi:\mathrm{Gal}(K(X[\ell])/K)\to\mathrm{End}^{0}(Y)^{*},$
whose image
$\Gamma:=\mathrm{Im}(\chi)\subset\mathrm{End}^{0}(Y)^{*}$
is a finite group that enjoys the following property.
The integers $n$ and #$\#(\Gamma)$ are not relatively prime. In other words, there is a prime $r\neq\ell$ that divides both $n$ and #$\#(\Gamma)$. In particular, both endomorphism algebras $\mathrm{End}^{0}(J^{(f,\ell)})$ and $\mathrm{End}^{0}(Y)$ contain an invertible element of multiplicative order $r$.
Proof of Theorem 2.2 (modulo Lemma 3.6).
It follows from Lemma 3.2 that $f(x)$ remains irreducible over $K(Y[\ell])$. So, replacing $K$ by $K(Y[\ell])$, we may and will assume that $K(Y[\ell])=K$. Now it follows from Lemma 3.6) that there is a prime $r\neq\ell$ that divides $n$ and enjoys the following property.
Both endomorphism algebras $\mathrm{End}^{0}(J^{(f,\ell)})$ and $\mathrm{End}^{0}(Y)$ contain an invertible element of multiplicative order $r$. This proves Theorem 2.2. ∎
Proof of Lemma 3.6.
In light of the theorem of Silverberg (Remark 2.1(ii)), all endomorphisms of $Y$ are defined over $K$. Applying this theorem (see Remark 2.1(ii) above) to $X\times Y$, we conclude that all the homomorphisms from $X$ to $Y$ are defined over $K(X[\ell])$.
Let $\mu:X\to Y$ be an isogeny. Dividing, if necessary, $\mu$ by a suitable power of $\ell$, we may and will assume that
$\mu(X[\ell])\neq\{0\}.$ (5)
Let us put
$G_{\ell}:=\tilde{G}_{\ell,X}=\mathrm{Gal}(K(X[\ell])/K),\ G=\mathrm{Gal}(K(X[1-\delta_{\ell}])/K)=\mathrm{Gal}(f/K).$
We know that $\mu$ is defined over $K(X[\ell])$. This allows us to define for each $\sigma\in G_{\ell}$ the isogeny $\sigma(\mu):X\to Y$, which is the Galois-conjugate of $\mu$ (recall that both $X$ and $Y$ are defined over $K$). Then the same construction as in [20, Sect. 4, proof of Prop. 2.4] allows us to define a map
$c:G_{\ell}\to\mathrm{End}^{0}(Y)^{*},\ \sigma\mapsto c(\sigma)$
where $c(\sigma)$ is determined by
$\sigma(\mu)=c(\sigma)\mu\ \forall\sigma\in G_{\ell}=\mathrm{Gal}(K(X[\ell])/K).$
We have for each $\sigma,\tau\in G_{\ell}$
$c(\sigma\tau)\mu=\sigma\tau(\mu)=\sigma(\tau(\mu))=\sigma(c(\tau)\mu)=c(\tau)\sigma(\mu)=c(\tau)c(\sigma)\mu$
(here we use that all elements of $\mathrm{End}(Y)$ are defined over $K$, i.e., are $G_{\ell}$-invariant). Therefore
$c(\sigma\tau)=c(\tau)c(\sigma)\ \forall\sigma,\tau\in G_{\ell}=\mathrm{Gal}(K(X[\ell])/K).$
This means that the map
$\chi:G_{\ell}=\mathrm{Gal}(K(X[\ell])/K)\to\mathrm{End}^{0}(Y)^{*},\ \sigma\mapsto\chi(\sigma)=c(\sigma)^{-1}$
is a group homomorphism. Let $\Gamma\subset\mathrm{End}^{0}(Y)^{*}$ be the image of $\chi$, which is a finite subgroup of $\mathrm{End}^{0}(Y)^{*}$. We need to check that there is a prime divisor $r$ of $n$ that divides #$\#(\Gamma)$.
Let $H_{\ell}$ be the kernel of $\chi$, i.e.,
$H_{\ell}=\{\sigma\in G_{\ell}\mid\sigma(\mu)=\mu\}.$ (6)
By definition, $H_{\ell}$ is a normal subgroup of $G_{\ell}$. Let $H$ be the image of $H_{\ell}$ in $G$ under the natural surjective group homomorphism
$G_{\ell}=\mathrm{Gal}(K(X[\ell])/K)\twoheadrightarrow\mathrm{Gal}(K(X[1-\delta_{\ell}])/K)=\mathrm{Gal}(f/K)=G$
induced by the inclusion
$K(X[1-\delta_{\ell}])\subset K(X[\ell])$
of Galois extensions of $K$. The surjectiveness implies that $H$ is a normal subgroup of $G$ and the index $(G:H)$ divides
$(G_{\ell}:H_{\ell})=\#(\Gamma).$
In order to finish the proof, we need the following assertion that will be proven at the end of this section.
Proposition 3.7.
The subgroup $H$ of $G$ is not transitive on $\mathfrak{R}_{f}$.
End of Proof of Lemma 3.6 (modulo Proposition 3.7) Combining Proposition 3.7 with Lemma 3.1, we conclude that $(G:H)$ is not prime to $n$. Hence, there is a prime $r$ that divides both $(G:H)$ and $n$. Since $n$ is prime to $\ell$ and $(G:H)$ divides $(G_{\ell}:H_{\ell})$, we conclude that $r\neq\ell$ and $r$ divides $(G_{\ell}:H_{\ell})=\#(\Gamma)$. This ends the proof.
Proof of Proposition 3.7.
Suppose that $H$ is transitive. Then $f(x)$ remains irreducible over the subfield $E:=K(\mathfrak{R}_{f})^{H}$ of all $H$-invariants in $K(\mathfrak{R}_{f}):=F$. Clearly, $E/K$ is a Galois extension and
$K\subset E\subset F=K(\mathfrak{R}_{f})=E(\mathfrak{R}_{f}).$
Let us consider the subfield $L:=K(X[\ell])^{H_{\ell}}$ of all $H_{\ell}$-invariants in $K(X[\ell])$, which is also a Galois extension of $K$. We have
$K(X[\ell])=L(X[\ell]),\quad H_{\ell}=\mathrm{Gal}(K(X[\ell])/L)=\mathrm{Gal}(L(X[\ell])/L).$
In addition,
$K\subset L=K(X[\ell])^{H_{\ell}}\supset K(\mathfrak{R}_{f})^{H}=F^{H}=E\supset K$
and
$E=F^{H}=F\cap K(X[\ell])^{H_{\ell}}=F\cap L,$
becase $H$ is the image of $H_{\ell}\subset\mathrm{Gal}(K(X[\ell])/K)$ in $\mathrm{Gal}(K(\mathfrak{R}_{f})/K)=\mathrm{Gal}(F/K)$.
We want to prove that $f(x)$ remains irreducible over $L$. In order to do it, notice that $H_{\ell}$ acts transitively on $\mathfrak{R}_{f}$. Indeed, let $\alpha,\beta$ be elements of $\mathfrak{R}_{f}$. By our assumption, there is $\sigma\in H\in\mathrm{Gal}(K(\mathfrak{R}_{f})/K)$ such that $\sigma(\alpha)=\beta$. Pick a field automorphism
$\sigma_{\ell}\in H_{\ell}\subset\mathrm{Gal}(K(X[\ell])/L)\subset\mathrm{Gal}(K(X[\ell])/K)$
such that the restriction of $\sigma_{\ell}$ to $K(\mathfrak{R}_{f})$ coincides with $\sigma$. Since
$\alpha\in\mathfrak{R}_{f}\subset K(\mathfrak{R}_{f}),$
we get
$\sigma_{\ell}(\alpha)=\sigma(\alpha)=\beta.$
This proves the transitivity of of the action of $H_{\ell}$ on the set $\mathfrak{R}_{f}$ of roots of $f(x)$. It follows that $f(x)$ is irreducible over the field $K(X[\ell])^{H_{\ell}}=L$.
Replacing $K$ by its overfield $L=K(X[\ell])^{H_{\ell}}$, we may and will assume that
$H_{\ell}=\mathrm{Gal}(K(X[\ell])/K).$
In particular,
$\sigma(\mu)=\mu\ \forall\sigma\in H_{\ell}=\mathrm{Gal}(K(X[\ell])/K).$ (7)
Recall that
$\sigma(\mu)(\sigma(x))=\sigma(\mu(x))\ \forall\sigma\in\mathrm{Gal}(K(X[\ell])/K),\ x\in X(K[\ell]).$ (8)
Since $X[\ell]\subset X(K[\ell])$ and nonzero $\mu(X[\ell])$ obviously lies in $Y[\ell]$, we conclude that the map
$X[\ell]\to Y[\ell],\ x\mapsto\mu(x)$ (9)
is a nonzero homomorphism of $\mathrm{Gal}(K)$-modules. Recall that we assume that the Galois module $Y[\ell]$ is trivial. On the other hand, in light of Corollary 3.5, the Galois module $X[\ell]$ admits a filtration, all whose consecutuve quotients are isomorphic to $Q_{\mathfrak{R}_{f}}$. Since $f(x)$ is irreducible over $K$, it follows from Lemma 2.4 that the homomorphism (9) is zero, which is not the case. The obtained contradiction proves that $H$ is not transitive. ∎
4. Proof of Theorem 1.1
So, $n$ is an odd prime, both $f(x)$ and $h(x)\in K[x]$ are degree $n$ polynomials without repeated roots, $f(x)$ is irreducible and $h(x)$ is reducible. Since $n$ is a prime, the reducibility of $h(x)$ implies that the order of $\mathrm{Gal}(h/K)$ is prime to $n$ (see [20, Lemma 2.6]). Let us put $Y=J^{(h,\ell)}$. We are given that the degree $[K(\mathfrak{R}_{h}):K]$ of the field extension $K(\mathfrak{R}_{h})/K$ is not divisible by $n$. Applying Corollary 3.5 to $Y$ and $h(x)$ (instead of $X=J^{(f,\ell)}$ and $f(x)$), we conclude that the degree of the field extension $K(Y[\ell])/K(\mathfrak{R}_{h})$ is either $1$ or a power of $\ell$. This implies that the order of the group $\tilde{G}_{\ell,Y}$ (which is
$[K(Y[\ell]):K]=[K(Y[\ell]):K(\mathfrak{R}_{h})]\cdot[K(\mathfrak{R}_{h}):K]\ \text{)}$
is prime to $n$, because the prime $\ell$ does not divide $n$. Now the desired result follows readily from Theorem 2.2, because if $r$ is a prime divisor of $n$ then $r=n$.